How much pure acid should be mixed with 6 gallons if a 60% acid in order to get a 70% acid solution
Answer by mananth(16946)

.......... Percent ---------------- quantity
Solution ..........60----------------30
Pure acid.........100----------------x
Total 70 ---------------- 30+x
60*6+100 *x=70(6+ x)
360+100x =420+70x
100x-70 x=420-360

30x=60
/30
x=2 gallons 100 % Pure acid